# Generalization under Byzantine & Poisoning Attacks: Tight Stability Bounds in Robust Distributed Learning

### v1

@@ D.1 Expectation bound for the honest gradients’ empirical covariance trace and spectral norm
[…]
    and
[…]
    Substituting our derivation into the bound from Lemma D.4 yields:
    $\mathbb{E}_{\mathcal{A}}[|\ell(\theta_{T};z)-\ell(\theta_{T}^{\prime};z)|]\leq\frac{t_{0}\ell_{\infty}}{m}+\frac{\sigma^{F}_{\mathrm{SGD}}C}{L}{\left(\frac{T}{t_{0}}\right)}^{c}$
    We approximately minimize the expression with respect to $t_{0}$ by balancing the two terms, i.e., setting them equal:
    $0\leq\widetilde{t_{0}}={\left(\frac{m\sigma^{F}_{\mathrm{SGD}}C}{L\ell_{\infty}}\right)}^{\frac{1}{c+1}}T^{\frac{c}{c+1}}\underset{\textit{for sufficiently large T}\geq\frac{m\sigma^{F}_{\mathrm{SGD}}C}{L\ell_{\infty}}}{\leq}T.$
    Finally, we obtain the result by directly substituting $\widetilde{t_{0}}$ into the bound:
    $\mathbb{E}_{\mathcal{A}}[|\ell(\theta_{T};z)-\ell(\theta_{T}^{\prime};z)|]\leq 2{\left(\frac{\sigma^{F}_{\mathrm{SGD}}C}{L}\right)}^{\frac{1}{c+1}}{\left(\frac{\ell_{\infty}T}{m}\right)}^{\frac{c}{c+1}}.\qed$
D.3 CWTM aggregation rule properties
We begin by analyzing the Lipschitz continuity of the Trimmed Mean ($\mathrm{TM}$) robust aggregation on the real line, which is useful for showing the expansivity property of the Component-Wise Trimmed Mean ($\mathrm{CWTM}$) robust update.
Lemma D.6.
The trimmed mean operation is $\frac{1}{n-2f}$-Lipschitz continuous with respect to the $\|\cdot\|_{1}$ norm, and $\frac{\sqrt{n}}{n-2f}$-Lipschitz continuous with respect to the $\|\cdot\|_{2}$ norm.
Proof.
We recall the definition of the trimmed mean operation: given $f,n\in\mathbb{N}$, $f<n/2$, and real values $x_{1},\ldots,x_{n}$, we denote by $\sigma$ a permutation on $\{1,\ldots,n\}$ that sorts the values in non-decreasing order $x_{\sigma(1)}\leq\cdots\leq x_{\sigma(n)}$. Let $S_{x}=\{i\in[n];f+1\leq\sigma(i)\leq n-f\}$, the trimmed mean is given by:
$TM(x_{1},\ldots,x_{n})=\frac{1}{n-2f}\sum_{i=f+1}^{n-f}x_{\sigma(i)}=\frac{1}{n-2f}\sum_{i\in S_{x}}x_{i}$
We first consider two sets of real numbers that differ in exactly one entry $\iota\in\{1,\ldots,n\}$: $x_{1},\ldots,x_{\iota},\ldots,x_{n}$ and $x^{\prime}_{1},\ldots,x^{\prime}_{\iota},\ldots,x^{\prime}_{n}$ such that $x_{j}=x^{\prime}_{j}$ for all $j\neq\iota$. Let $S_{x}$ and $S_{x^{\prime}}$ denote the sets of selected indices in each case. We necessarily have $|S_{x^{\prime}}\setminus S_{x}|\leq 1$. In the figure below, we illustrate different scenarios to provide an intuition for our proof.
$x_{1}$$x_{2}$$\underline{\iota}$$x_{3}$$x_{4}$$x_{5}$$\overline{\iota}$$x_{6}$($\mathrm{TM}(x)$$x_{4}\neq x_{4}^{\prime}$$x_{1}$$x_{2}$$x_{3}$$x_{4}^{\prime}$$x_{5}$$x_{6}$($\mathrm{TM}(x^{\prime})$$x_{6}\neq x_{6}^{\prime}$$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$$x_{6}^{\prime}$($\mathrm{TM}(x^{\prime})$
Figure 2: Example with $n=6$, $f=2$ and the differing point being either $x_{6}$ or $x_{4}$. Horizontal dashed lines represent either trimmed mean differences or changed-point differences. Green points indicate the contributors to the trimmed-mean, while red points denote those excluded from the computation.
We consider the two possible cases.
  • Case (i) $|S_{x^{\prime}}\setminus S_{x}|=1$. Let $x_{\underline{\iota}}$ and $x_{\overline{\iota}}$ be the boundary elements not included in $S_{x}$. That is, if we denote by $\sigma$ a permutation on $\{1,\ldots,n\}$ that sorts the values in non-decreasing order $x_{\sigma(1)}\leq\cdots\leq x_{\sigma(n)}$, then $\underline{\iota}$ and $\overline{\iota}$ are defined as follow: $\sigma(\underline{\iota})=f$ and $\sigma(\overline{\iota})=n-f+1$.
    • $\circ$
      If $\iota\notin S_{x}$ and $\iota\notin S_{x^{\prime}}$, then with $\{j\}=S_{x}\setminus S_{x^{\prime}}$, we have two sub-cases:
      • Sub-case (a) $x_{\iota}>x_{\overline{\iota}}$ and $x_{\iota}^{\prime}<x_{\underline{\iota}}$. In this case, $S_{x^{\prime}}\setminus S_{x}=\{x_{\underline{\iota}}\}$. Therefore,
        $\textstyle|\mathrm{TM}(x)-\mathrm{TM}(x^{\prime})|=\frac{1}{n-2f}|x_{j}-x_{\underline{\iota}}|\leq\frac{1}{n-2f}|x_{\overline{\iota}}-x_{\underline{\iota}}|\leq\frac{1}{n-2f}|x_{\iota}-x_{\iota}^{\prime}|.$
      • Sub-case (b) $x_{\iota}<x_{\underline{\iota}}$ and $x_{\iota}^{\prime}>x_{\overline{\iota}}$. In this case, $S_{x^{\prime}}\setminus S_{x}=\{x_{\overline{\iota}}\}$. Therefore,
        $\textstyle|\mathrm{TM}(x)-\mathrm{TM}(x^{\prime})|=\frac{1}{n-2f}|x_{\overline{\iota}}-x_{j}|\leq\frac{1}{n-2f}|x_{\overline{\iota}}-x_{\underline{\iota}}|\leq\frac{1}{n-2f}|x_{\iota}-x_{\iota}^{\prime}|.$
    • $\circ$
      If $\iota\in S_{x}$ but $\iota\notin S_{x^{\prime}}$ then we have the following two sub-cases:
      • Sub-case (c) $x_{\underline{\iota}}\in S_{x^{\prime}}$. In this case, $x_{\iota}^{\prime}\leq x_{\underline{\iota}}$. Therefore,
        $\textstyle|\mathrm{TM}(x)-\mathrm{TM}(x^{\prime})|=\frac{1}{n-2f}|x_{\underline{\iota}}-x_{\iota}|\leq\frac{1}{n-2f}|x^{\prime}_{\iota}-x_{\iota}|.$
      • Sub-case (d) $x_{\overline{\iota}}\in S_{x^{\prime}}$. In this case, $x_{\iota}^{\prime}\geq x_{\overline{\iota}}$. Therefore,
        $\textstyle|\mathrm{TM}(x)-\mathrm{TM}(x^{\prime})|=\frac{1}{n-2f}|x_{\overline{\iota}}-x_{\iota}|\leq\frac{1}{n-2f}|x^{\prime}_{\iota}-x_{\iota}|.$
    • $\circ$
      By symmetry, we obtain the same result for the case when $\iota\notin S_{x}$ but $\iota\in S_{x^{\prime}}$.
  • Case (ii) $|S_{x^{\prime}}\setminus S_{x}|=0$. The lemma is trivially true in this case.
From above, we obtain that $|\mathrm{TM}(x)-\mathrm{TM}(x^{\prime})|\leq\frac{1}{n-2f}\|x-x^{\prime}\|_{1}$.
We can generalize this bound to two sets of $n$ real numbers that can differ in all the entries, using triangle inequality. Specifically,
$|\mathrm{TM}(x)-\mathrm{TM}(x^{\prime})|$ $=|\mathrm{TM}(x)-\mathrm{TM}(x_{1},\ldots,x_{n-1},x^{\prime}_{n})$
$\qquad+\mathrm{TM}(x_{1},\ldots,x_{n-1},x^{\prime}_{n})-\mathrm{TM}(x_{1},\ldots,x_{n-2},x^{\prime}_{n-1},x^{\prime}_{n})$
$\qquad+\cdots+\mathrm{TM}(x_{1},x^{\prime}_{2},\ldots,x^{\prime}_{n})-\mathrm{TM}(x^{\prime})|$
$\leq|\mathrm{TM}(x)-\mathrm{TM}(x_{1},\ldots,x_{n-1},x^{\prime}_{n})|$
$\qquad+\cdots+|\mathrm{TM}(x_{1},x^{\prime}_{2},\ldots,x^{\prime}_{n})-\mathrm{TM}(x^{\prime})|$
$\leq\frac{1}{n-2f}\sum_{i=1}^{n}|x_{i}-x^{\prime}_{i}|=\frac{1}{n-2f}\|x-x^{\prime}\|_{1}\leq\frac{\sqrt{n}}{n-2f}\|x-x^{\prime}\|_{2}\qed$
We next present another property of the Trimmed Mean that, like Lipschitz continuity, enables us to establish the expansivity of the Component-Wise Trimmed Mean update rule.
Lemma D.7.
Let $x,x^{\prime}\in\mathbb{R}^{n}$. There exist indices $\alpha,\beta\in\{1,\ldots,n\}$ such that:
${\left[x-x^{\prime}\right]}_{\beta}\leq\mathrm{TM}(x)-\mathrm{TM}(x^{\prime})\leq{\left[x-x^{\prime}\right]}_{\alpha}.$
Proof.
Let $n\in\mathbb{N}$, $f\in\{0,\ldots,\lfloor\frac{n}{2}\rfloor\}$, and $x,x^{\prime}\in\mathbb{R}^{n}$
  • (i)
    Suppose $\mathrm{TM}(x)=\mathrm{TM}(x^{\prime})$. We then consider two cases:
    • If there exists an index $i\in\{1,\ldots,n\}$ such that $x_{i}=x^{\prime}_{i}$, the lemma is trivially true with $\alpha=\beta=i$.
    • Otherwise, $x_{i}\neq x^{\prime}_{i},\forall i\in\{1,\ldots,n\}$. In this case, there must be at least one index $\beta$ where ${\left[x-x^{\prime}\right]}_{\beta}<0$ and at least one index $\alpha$ where ${\left[x-x^{\prime}\right]}_{\alpha}>0$. If all differences were strictly positive or strictly negative, this would contradict the assumption that $\mathrm{TM}(x)=\mathrm{TM}(x^{\prime})$.
  • (ii)
    Suppose $\mathrm{TM}(x)\neq\mathrm{TM}(x^{\prime})$. Here, we invoke the translation equivariance property of the trimmed mean. Specifically, let $\mathbf{1}=(1,\ldots,1)\in\mathbb{R}^{n}$, and define the translation operator $\mathcal{T}_{a\mathbf{1}}(x)\vcentcolon=x-a\mathbf{1}$ for any $a\in\mathbb{R}$. Then, for any $a\in\mathbb{R}$, the trimmed mean satisfies:
    $\mathrm{TM}(\mathcal{T}_{a\mathbf{1}}(x))=\mathrm{TM}(x-a\mathbf{1})=\mathrm{TM}(x)-a$
    Substituting $a=\mathrm{TM}(x)-\mathrm{TM}(x^{\prime})$, we have $\mathrm{TM}(\mathcal{T}_{a\mathbf{1}}(x))-\mathrm{TM}(x^{\prime})=0$. We then invoke the first case: there exist indices $\alpha,\beta\in\{1,\ldots,n\}$ such that:
    ${\left[\mathcal{T}_{a\mathbf{1}}(x)-x^{\prime}\right]}_{\beta}\leq\mathrm{TM}(\mathcal{T}_{a\mathbf{1}}(x))-\mathrm{TM}(x^{\prime})\leq{\left[\mathcal{T}_{a\mathbf{1}}(x)-x^{\prime}\right]}_{\alpha},$
    $\implies{\left[x-x^{\prime}\right]}_{\beta}-a\leq\mathrm{TM}(x)-\mathrm{TM}(x^{\prime})-a\leq{\left[x-x^{\prime}\right]}_{\alpha}-a,$
    $\implies{\left[x-x^{\prime}\right]}_{\beta}\leq\mathrm{TM}(x)-\mathrm{TM}(x^{\prime})\leq{\left[x-x^{\prime}\right]}_{\alpha}.$
    This concludes the proof.∎
Lemma D.6 and Lemma D.7 allow us to establish the expansivity of the Component-Wise Trimmed Mean update in the case of nonconvex loss functions, as shown in the following result.
Lemma D.8 (Expansivity of robust stochastic gradient update with $\mathrm{CWTM}$).
Define the robust gradient update with learning rate $\gamma>0$, for arbitrary $z^{(i)}\in\mathcal{Z}$, any $i\in\{1,\ldots,n\}$, and any coordinate $k\in\{1,\ldots,d\}$, as follows:
${\left[G^{\mathrm{CWTM}}_{\gamma}(\theta)\right]}_{k}={\left[\theta\right]}_{k}-\gamma\mathrm{TM}\left({\left[\nabla\ell(\theta;z^{(1)})\right]}_{k},\ldots,{\left[\nabla\ell(\theta;z^{(n)})\right]}_{k}\right)\\ ={\left[\theta\right]}_{k}-\frac{\gamma}{n-2f}\sum_{i=f+1}^{n-f}{\left[\nabla\ell(\theta;z^{(\sigma_{k}(i))})\right]}_{k}={\left[\theta\right]}_{k}-\frac{\gamma}{n-2f}\sum_{i\in S^{(k)}}{\left[\nabla\ell(\theta;z^{(i)})\right]}_{k}.$
where $\sigma_{k}$ denotes the permutation that sorts the values $({\left[\nabla\ell(\theta;z^{(1)})\right]}_{k},\ldots,{\left[\nabla\ell(\theta;z^{(n)})\right]}_{k})$ in ascending order, and $S^{(k)}=\{i\in[n];f+1\leq\sigma_{k}(i)\leq n-f\}$ denotes the index set of central values in coordinate $k$. We can write,
$G^{\mathrm{CWTM}}_{\gamma}(\theta)=\theta-\gamma\mathrm{CWTM}(\nabla\ell(\theta;z^{(1)}),\ldots,\nabla\ell(\theta;z^{(n)})).$
Assume $\forall z\in\mathcal{Z},\ell(\cdot,z):\Theta\to\mathbb{R}$ is L-smooth and nonconvex. Then $G^{\mathrm{CWTM}}_{\gamma}$ is $(1+\gamma L\min\{\frac{n}{n-2f},\sqrt{n},\sqrt{d}\})$-expansive.
Proof.
Let $\theta,\omega\in\mathbb{R}^{d}$ and $z^{(1)},\ldots,z^{(n)}\in\mathcal{Z}$. For convenience, we occasionally adopt the shorthand notation: $\ell(\cdot;z^{(i)})=\ell_{i}(\cdot)$. Using Lemma D.6 and the $L$-smoothness assumption, we obtain that
$\|\mathrm{CWTM}(\nabla\ell_{1}(\theta),\ldots,\nabla\ell_{n}(\theta))-\mathrm{CWTM}(\nabla\ell_{1}(\omega),\ldots,\nabla\ell_{n}(\omega))\|_{2}^{2}$
$=\sum_{k=1}^{d}{\left(\mathrm{TM}({\left[\nabla\ell_{1}(\theta)\right]}_{k},\ldots,{\left[\nabla\ell_{n}(\theta)\right]}_{k})-\mathrm{TM}({\left[\nabla\ell_{1}(\omega)\right]}_{k},\ldots,{\left[\nabla\ell_{n}(\omega)\right]}_{k})\right)}^{2}$
$\leq\frac{n}{{(n-2f)}^{2}}\sum_{k=1}^{d}\|({\left[\nabla\ell_{1}(\theta)\right]}_{k},\ldots,{\left[\nabla\ell_{n}(\theta)\right]}_{k})-({\left[\nabla\ell_{1}(\omega)\right]}_{k},\ldots,{\left[\nabla\ell_{n}(\omega)\right]}_{k})\|_{2}^{2}$
$=\frac{n}{{(n-2f)}^{2}}\sum_{k=1}^{d}\sum_{i=1}^{n}{\left({\left[\nabla\ell_{i}(\theta)\right]}_{k}-{\left[\nabla\ell_{i}(\omega)\right]}_{k}\right)}^{2}=\frac{n}{{(n-2f)}^{2}}\sum_{i=1}^{n}\|\nabla\ell_{i}(\theta)-\nabla\ell_{i}(\omega)\|_{2}^{2}$
$\leq\frac{n}{{(n-2f)}^{2}}\sum_{i=1}^{n}L^{2}\|\theta-\omega\|_{2}^{2}=\frac{n^{2}L^{2}}{{(n-2f)}^{2}}\|\theta-\omega\|_{2}^{2}.$ (11)
Alternatively, we can invoke Lemma D.7, which states that $\forall k\in\{1,\ldots,d\}$, there exists $i_{k}\in\{1,\ldots,n\}$ such that:
$\|\mathrm{CWTM}(\nabla\ell_{1}(\theta),\ldots,\nabla\ell_{n}(\theta))-\mathrm{CWTM}(\nabla\ell_{1}(\omega),\ldots,\nabla\ell_{n}(\omega))\|_{2}^{2}$
$=\sum_{k=1}^{d}{\left(\mathrm{TM}({\left[\nabla\ell_{1}(\theta)\right]}_{k},\ldots,{\left[\nabla\ell_{n}(\theta)\right]}_{k})-\mathrm{TM}({\left[\nabla\ell_{1}(\omega)\right]}_{k},\ldots,{\left[\nabla\ell_{n}(\omega)\right]}_{k})\right)}^{2}$
$\leq\sum_{k=1}^{d}{\left[\nabla\ell_{i_{k}}(\theta)-\nabla\ell_{i_{k}}(\omega)\right]}_{k}^{2}.$
From above, we obtain the following:
  1. 1.
    Since $\sum_{k=1}^{d}{\left[\nabla\ell_{i_{k}}(\theta)-\nabla\ell_{i_{k}}(\omega)\right]}_{k}^{2}\leq\sum_{k=1}^{d}\sum_{i=1}^{n}{\left[\nabla\ell_{i}(\theta)-\nabla\ell_{i}(\omega)\right]}_{k}^{2}\leq nL^{2}\|\theta-\omega\|_{2}^{2}$ ,
    $\|\mathrm{CWTM}(\nabla\ell_{1}(\theta),\ldots,\nabla\ell_{n}(\theta))-\mathrm{CWTM}(\nabla\ell_{1}(\omega),\ldots,\nabla\ell_{n}(\omega))\|_{2}^{2}\leq nL^{2}\|\theta-\omega\|_{2}^{2}.$
  2. 2.
    Since $\sum_{k=1}^{d}{\left[\nabla\ell_{i_{k}}(\theta)-\nabla\ell_{i_{k}}(\omega)\right]}_{k}^{2}\leq\sum_{k=1}^{d}\|\nabla\ell_{i_{k}}(\theta)-\nabla\ell_{i_{k}}(\omega)\|^{2}_{2}$,
    $\|\mathrm{CWTM}(\nabla\ell_{1}(\theta),\ldots,\nabla\ell_{n}(\theta))-\mathrm{CWTM}(\nabla\ell_{1}(\omega),\ldots,\nabla\ell_{n}(\omega))\|_{2}^{2}\leq dL^{2}\|\theta-\omega\|_{2}^{2}.$
Therefore,
$\|\mathrm{CWTM}(\nabla\ell_{1}(\theta),\ldots,\nabla\ell_{n}(\theta))-\mathrm{CWTM}(\nabla\ell_{1}(\omega),\ldots,\nabla\ell_{n}(\omega))\|_{2}^{2}\leq\min\{n,d\}L^{2}\|\theta-\omega\|_{2}^{2}.$ (12)
The bound in (12) has a tighter dependence on the number of workers than the bound in (11), especially in the regime when $f$ approaches $\frac{n}{2}$. We conclude the proof by combining (11) and (12), and by applying the triangle inequality, yielding:
$\|G^{\mathrm{CWTM}}_{\gamma}(\theta)-G^{\mathrm{CWTM}}_{\gamma}(\omega)\|_{2}\leq\left(1+\gamma L\min\left\{\frac{n}{n-2f},\sqrt{n},\sqrt{d}\right\}\right)\|\theta-\omega\|_{2}.\qed$
This behavior contrasts with the convex and strongly convex settings, where, as illustrated below in Remark D.3, a non-expansive update (i.e., with expansivity coefficient $\leq 1$) cannot, in general, be guaranteed. Notably, the non-expansivity property is retained in one-dimensional settings or when the loss function decomposes across dimensions, akin to the behavior observed with averaging-based updates. This observation is formalized in the following lemma.
Lemma D.9.
Let $d=1$. Assume that for all $z\in\mathcal{Z}$, the loss function $\ell(\cdot;z)$ is convex and $L$-smooth. If $\gamma\leq\frac{2}{L}$ then $G^{\mathrm{CWTM}}_{\gamma}$ is $1$-expansive.
Proof.
We use the same notation as in Lemma D.8, with $\ell^{\prime}(\cdot)$ denoting the derivative of $\ell(\cdot)$. We obtain
$|G^{\mathrm{CWTM}}_{\gamma}(\theta)-G^{\mathrm{CWTM}}_{\gamma}(\omega)|^{2}=$ $|\theta-\omega|^{2}+\gamma^{2}|\mathrm{TM}(\ell^{\prime}_{1}(\theta),\ldots,\ell^{\prime}_{n}(\theta))-\mathrm{TM}(\ell^{\prime}_{1}(\omega),\ldots,\ell^{\prime}_{n}(\omega))|^{2}$
$-2\gamma(\theta-\omega)(\mathrm{TM}(\ell^{\prime}_{1}(\theta),\ldots,\ell^{\prime}_{n}(\theta))-\mathrm{TM}(\ell^{\prime}_{1}(\omega),\ldots,\ell^{\prime}_{n}(\omega))).$ (13)
In this case, we can leverage the co-coercivity property of the loss function. In fact, due to Lemma D.7, there exist indices $\alpha,\beta\in\{1,\ldots,n\}$ such that:
$\ell^{\prime}_{\beta}(\theta)-\ell^{\prime}_{\beta}(\omega)\leq\mathrm{TM}(\ell^{\prime}_{1}(\theta),\ldots,\ell^{\prime}_{n}(\theta))-\mathrm{TM}(\ell^{\prime}_{1}(\omega),\ldots,\ell^{\prime}_{n}(\omega))\leq\ell^{\prime}_{\alpha}(\theta)-\ell^{\prime}_{\alpha}(\omega).$ (14)
Therefore, there exists $i\in\{\alpha,\beta\}$ such that
$(\theta-\omega)\left(\mathrm{TM}(\ell^{\prime}_{1}(\theta),\ldots,\ell^{\prime}_{n}(\theta))-\mathrm{TM}(\ell^{\prime}_{1}(\omega),\ldots,\ell^{\prime}_{n}(\omega))\right)$ $\geq(\theta-\omega)(\ell^{\prime}_{i}(\theta)-\ell^{\prime}_{i}(\omega))$
$\geq\frac{1}{L}{(\ell^{\prime}_{i}(\theta)-\ell^{\prime}_{i}(\omega))}^{2}\geq 0.$
Thus, we have
$(\theta-\omega)(\mathrm{TM}(\ell^{\prime}_{1}(\theta),\ldots,\ell^{\prime}_{n}(\theta))-\mathrm{TM}(\ell^{\prime}_{1}(\omega),\ldots,\ell^{\prime}_{n}(\omega)))$
$=|\theta-\omega|\,|\mathrm{TM}(\ell^{\prime}_{1}(\theta),\ldots,\ell^{\prime}_{n}(\theta))-\mathrm{TM}(\ell^{\prime}_{1}(\omega),\ldots,\ell^{\prime}_{n}(\omega))|.$
Using this in (13), we obtain that
$|G^{\mathrm{CWTM}}_{\gamma}(\theta)-G^{\mathrm{CWTM}}_{\gamma}(\omega)|^{2}=$ $|\theta-\omega|^{2}+\gamma^{2}|\mathrm{TM}(\ell^{\prime}_{1}(\theta),\ldots,\ell^{\prime}_{n}(\theta))-\mathrm{TM}(\ell^{\prime}_{1}(\omega),\ldots,\ell^{\prime}_{n}(\omega))|^{2}$
$-2\gamma|\theta-\omega|\,|\mathrm{TM}(\ell^{\prime}_{1}(\theta),\ldots,\ell^{\prime}_{n}(\theta))-\mathrm{TM}(\ell^{\prime}_{1}(\omega),\ldots,\ell^{\prime}_{n}(\omega))|.$ (15)
Note that, due to (14), there also exists $j\in\{\alpha,\beta\}$ such that
$|\mathrm{TM}(\ell^{\prime}_{1}(\theta),\ldots,\ell^{\prime}_{n}(\theta))-\mathrm{TM}(\ell^{\prime}_{1}(\omega),\ldots,\ell^{\prime}_{n}(\omega))|\leq|\ell^{\prime}_{j}(\theta)-\ell^{\prime}_{j}(\omega)|\leq L|\theta-\omega|.$ (16)
Substituting from (16) in (15), we obtain that
$|G^{\mathrm{CWTM}}_{\gamma}(\theta)-G^{\mathrm{CWTM}}_{\gamma}(\omega)|^{2}\leq|\theta-\omega|^{2}$
$+\gamma(\gamma L-2)|\theta-\omega|\,|\mathrm{TM}(\ell^{\prime}_{1}(\theta),\ldots,\ell^{\prime}_{n}(\theta))-\mathrm{TM}(\ell^{\prime}_{1}(\omega),\ldots,\ell^{\prime}_{n}(\omega))|.$
The above proves the lemma since $(\gamma L-2)\leq 0$ when $\gamma\leq\frac{2}{L}$. ∎
Remark D.2.
Lemma D.9 generalizes to loss functions that separate across dimensions; that is, functions of the form $\ell=\sum_{k=1}^{d}\ell_{k}$, where each $\ell_{k}$ is a one dimensional function acting in the $k$-th coordinate, $k\in\{1,\ldots,d\}$. A canonical example is mean estimation under coordinate-wise squared error loss, where the total loss decomposes as a sum of independent scalar losses for each dimension.
Remark D.3.
To establish, in general, a non-expansivity result for the $\mathrm{CWTM}$ update rule, it would be necessary that the scalar product in:
$\|G^{\mathrm{CWTM}}_{\gamma}(\theta)-G^{\mathrm{CWTM}}_{\gamma}(\omega)\|_{2}^{2}\\ =\|\theta-\omega\|_{2}^{2}+\gamma^{2}\|\mathrm{CWTM}(\nabla\ell_{1}(\theta),\ldots,\nabla\ell_{n}(\theta))-\mathrm{CWTM}(\nabla\ell_{1}(\omega),\ldots,\nabla\ell_{n}(\omega))\|_{2}^{2}\\ -2\gamma\langle\mathrm{CWTM}(\nabla\ell_{1}(\theta),\ldots,\nabla\ell_{n}(\theta))-\mathrm{CWTM}(\nabla\ell_{1}(\omega),\ldots,\nabla\ell_{n}(\omega)),\theta-\omega\rangle$
is strictly positive, that is the $\mathrm{CWTM}$ preserve the co-coercivity of the underlying loss function. As we will illustrate bellow, this is not the case. This is because the sets of selected indices differ across dimensions—that is $S^{(k)}\neq S^{(k)\prime}$ with $k\in\{1,\ldots,d\}$—which disrupts the alignment needed to preserve the co-coercivity property.
[…]

### v2

@@ C.1 Expectation bounds for honest gradients’ empirical spectral norm
  • $\text{Var}({\{g^{(i)}_{t}\}}_{i\in\mathcal{H}})=C^{2}$, and any subset $\text{Select}_{a}(\mathcal{H}_{-})\cup\text{Select}_{b}(\mathcal{H}_{+})\cup\text{Select}_{n-f-a-b}(\mathcal{B})$, with $0<a<\frac{n-f}{2}$, $0\leq b<\frac{n-f}{2}$ yields higher variance.
  • As $f\geq n/3$, $f\geq\frac{n-f}{2}\geq n-2f$. Hence, either $f=\frac{n-f}{2}$ and
    $\mathcal{H}_{+}\cup\text{Select}_{\frac{n-3f}{2}}(\mathcal{B})=\text{Select}_{n-2f}(\mathcal{H}_{+})\cup\mathcal{B}=\mathcal{H}_{+}\cup\mathcal{B},$
    or $f>\frac{n-f}{2}$ and $\text{Var}({\{g^{(i)}_{t}\}}_{i\in\mathcal{H}_{+}\cup\text{Select}_{\frac{n-3f}{2}}(\mathcal{B})})>\text{Var}({\{g^{(i)}_{t}\}}_{i\in\text{Select}_{n-2f}(\mathcal{H}_{+})\cup\mathcal{B}})$. Hence, the aggregation rule will always pick a larger subgroup over a smaller one (it minimizes the variance magnitude).
  • $\text{Var}({\{g^{(i)}_{t}\}}_{i\in\text{Select}_{n-2f}(\mathcal{H}_{+})\cup\mathcal{B}})=\frac{C^{2}}{4}<C^{2}=\text{Var}({\{g^{(i)}_{t}\}}_{i\in\mathcal{H}})$. In fact,
    $\text{Var}({\{g^{(i)}_{t}\}}_{i\in\text{Select}_{n-2f}(\mathcal{H}_{+})\cup\mathcal{B}})$ $=\frac{f}{n-f}{\left[\beta-\frac{1}{n-f}\left(f\beta+(n-2f)C\right)\right]}^{2}$
    $\qquad+\frac{n-2f}{n-f}{\left[C-\frac{1}{n-f}\left(f\beta+(n-2f)C\right)\right]}^{2}$
    $=\frac{f(n-2f)}{{(n-f)}^{2}}{\left(\beta-C\right)}^{2}$
    $=\frac{f(n-2f)}{{(n-f)}^{2}}C^{2}{\left(1+\frac{1}{2}\frac{n-f}{\sqrt{f(n-2f)}}-1\right)}^{2}$
    $=\frac{C^{2}}{4}.$
This reasoning applies to all iterations because the gradients, by construction, stay constant during the entire optimization process. As a result,
$\theta_{T}$ $=\theta_{0}-\gamma T\frac{1}{n-f}\sum_{i\in\text{Select}_{n-2f}(\mathcal{H}_{+})\cup\mathcal{B}}g^{(i)}_{0}$
$=-\frac{\gamma T}{n-f}\left(fC\left(1+\frac{1}{2}\frac{n-f}{\sqrt{f(n-2f)}}\right)+(n-2f)C\right)$
$=-\gamma TC\left(1+\frac{1}{2}\sqrt{\frac{f}{n-2f}}\right).$
Hence,
$\sup_{z\in[-C,C]}|\ell(\theta_{T};z)-\ell(\theta^{\prime}_{T};z)|$ $=\sup_{z\in[-C,C]}z\left(|\theta_{T}-\theta^{\prime}_{T}|\right)$
$=\gamma C^{2}T\left(1+\frac{1}{2}\sqrt{\frac{f}{n-2f}}+\frac{1}{(n-f)m}\right)\in\Omega\left(\gamma C^{2}T\left(\sqrt{\frac{f}{n-2f}}+\frac{1}{(n-f)m}\right)\right).$
(ii) If nfn-f is odd and $f\geq\frac{n}{3}$.
The same construction works with a slight adaptation. That is, we define the following pair of neighboring datasets:
$\begin{split}S=\{D_{1},\ldots,D_{n-f}\},&\quad S^{\prime}=S\setminus\{D_{1}\}\cup\{D_{1}\setminus\{z^{(1,1)}\}\cup\{z^{\prime(1,1)}\}\}\\ \text{ with }\forall i\in\{1,\ldots,n-f\},&\quad D_{i}=\{z^{(i,1)},\ldots,z^{(i,m)}\}\\ \text{ and }\forall j\in\{1,\ldots,m\},&\quad i\in\mathcal{H}_{+}=\{1,\ldots,\frac{n-f-1}{2}\},|\mathcal{H}_{+}|=\frac{n-f-1}{2}\!\!:z^{(i,j)}=C\text{ and }z^{\prime(1,1)}=0.\\ &\quad i\in\mathcal{H}_{-}=\{\frac{n-f-1}{2}+1,\ldots,n-f-1\},|\mathcal{H}_{-}|=\frac{n-f-1}{2}\!\!:z^{(i,j)}=-C.\\ &\quad z^{(n-f,j)}=0.\end{split}$
Then,
$\text{Var}({\{g^{(i)}_{t}\}}_{i\in\mathcal{H}})=(1-\frac{1}{n-f})C^{2}>\frac{C^{2}}{3}\text{ as }n-f>\frac{n}{2}\geq\frac{3}{2}.$
The remainder of the proof when nfn-f is odd proceeds identically to the previous case when nfn-f is even.
(iii) In the regime $f<n/3$.
For values restricted to $z\in[-C,C]$, variance is maximized by placing all points at the endpoints $\pm C$, with counts as balanced as possible. Thus we take
$k=\left\lceil\frac{n-f}{2}\right\rceil\quad\text{points at }+C,\qquad\left\lfloor\frac{n-f}{2}\right\rfloor=n-f-k\quad\text{points at }-C.$
We have two cases, if nfn-f is even then the variance equals $C^{2}$, else it equals $C^{2}\left(1-\frac{1}{{(n-f)}^{2}}\right)$. Because $f<n/3$ we have n2f>(nf)/2n-2f>(n-f)/2. Hence, we would like to consider the subset $S$ composed by $f$ times $\beta$, $k$ times $C$ and n2fkn-2f-k times C-C. We compute the competing set variance by the identity $\mathrm{Var}(X)=\mathbb{E}[X^{2}]-\mathbb{E}[X]^{2}$,
$\mathrm{Var}(S)=\frac{(n-2f)C^{2}+f\beta^{2}}{n-f}-\frac{\big((2k-n+2f)C+f\beta\big)^{2}}{(n-f)^{2}}.$
Requiring $\mathrm{Var}(S)\leq\mathrm{Var}(\mathcal{H})=C^{2}\!\left(1-\frac{(2k-n+f)^{2}}{(n-f)^{2}}\right)$ yields the quadratic inequality in $\beta$
$(n-2f)\beta^{2}+2(n-2f-2k)C\beta-(4k-n+2f)C^{2}\leq 0.$
The discriminant simplifies exactly to $16C^{2}k^{2}$, so the two roots are
$\beta_{-}=-C,\qquad\beta_{+}=\frac{4k-(n-2f)}{n-2f}C=(\frac{4k}{n-2f}-1)C.$
Since (n2f)>0(n-2f)>0, the admissible $\beta$ lie between the two roots. Moreover, since $\frac{2k}{n-2f}=\frac{2\left\lceil\frac{n-f}{2}\right\rceil}{n-2f}\geq\frac{n-f}{n-2f}$, we have $\frac{4k}{n-2f}-1\geq 2\frac{n-f}{n-2f}-1=\frac{2f}{n-2f}+1>1$.
Next, we have:
$\theta_{T}=-\gamma CT\frac{2k-n+2f+\frac{4fk}{n-2f}-f}{n-f}=-\gamma CT\left(\frac{4fk}{(n-f)(n-2f)}+\frac{2k-(n-f)}{n-f}\right),$
with $2k\geq n-f$. Hence, similarly to the previous cases, we have $\theta_{T}^{\prime}>0$, resulting in:
[…]